### v1

@@ 5.4 Coming Out versus LGBT Pride Month
As described in Section 2.2, the disclosure of Graves’ sexual orientation coincided with the start of LGBT Pride Month. This means that the coming-out event encompasses two “simultaneous treatments” [[, see, e.g.,]]roller2023differences, namely the announcement of Graves’ homosexuality and the introduction of visual and expressive elements in League of Legends that support the LGBT community. It is therefore plausible that the findings presented in Section 4.1 may, to some extent, be influenced by the presence of LGBT Pride Month, which might elicit negative reactions from certain players, leading them to shift their preferences away from LGB characters. While this alternative perspective does not undermine the validity of our identification strategy, it does raise questions about our interpretation of the estimated effects as solely stemming from Graves’ disclosure.
@@
Figure 5.4: Composite LGB unit’s daily pick rates and synthetic control estimation results. The actual series is smoothed by a Nadaraya-Watson regression before applying the synthetic control estimator. The dashed vertical line denotes the day of disclosure, and the rainbow area highlights LGBT Pride Month.
We create a composite LGB unit by averaging the pick rates of Diana, Leona, Nami, and Neeko and employ the synthetic control estimator described in Section 3.2 to estimate the effect of LGBT Pride Month on players’ preferences for LGB characters. Then, under the assumption that the influence of LGBT Pride Month is uniform across all LGB characters, we can compare the results with those obtained for Graves to separate the impacts of coming out and LGBT Pride Month on players’ preferences for Graves. Intuitively, if the estimated impact of LGBT Pride Month on players’ preferences for LGB characters is small relative to the estimated impact of the coming-out event on players’ preferences for Graves, this suggests that the findings of Section 4.1 must be primarily attributed to Graves’ disclosure.
Figure 5.4 displays the actual and the synthetic pick rate series for the composite LGB unit. Overall, our analysis suggests that LGBT Pride Month had no impact on players’ preferences for LGB characters. Before the treatment, the synthetic control estimator closely aligns with the actual series, providing support for the estimator’s ability to predict the counterfactual series. After the treatment date, the synthetic control estimator continues to align with the actual series, confirming that the players’ preferences for LGB characters were unaffected by LGBT Pride Month. The average effect is estimated to be $-0.149$ percentage points (standard error: $2.832$), and the conventional $95\%$ confidence interval encompasses zero, indicating a failure to reject the null hypothesis of no effect. Under the homogeneity assumption discussed above, these findings support the interpretation that the estimated effects presented in Section 4.1 are primarily driven by Graves’ disclosure rather than being influenced by the broader context of LGBT Pride Month.
@@ 6 Conclusion
To bolster the credibility of stigma as the primary explanation for the estimated effects, we address and eliminate several alternative channels. First, we rule out the possibility that shifts in characters’ relative strengths could explain our estimated effect. Second, we show that players’ skills have no correlation with the choice to drop the character. Third, we provide evidence that switching to other characters does not affect the performance of the players involved. Fourth, we introduce a theoretical framework that formalizes the existence of two “simultaneous treatments” and use information on other playable characters that belong to the LGB group to rule out that LGBT Pride Month serves as the main explanation for the observed result.
Our data and institutional setting offer great advantages, as they allow for a credible identification and reduced social desirability bias to identify attitudes towards LGB status [202]. However, it is important to consider at least three aspects when thinking about the implications of our results. First, our population is likely not representative of the general population, as the former is possibly less diverse. We recognize that responses might be different in a more diverse audience. Second, we study the responses of a male character revealing his sexual minority status. In line with previous evidence suggesting differential labor market discrimination for gay men and lesbian women, [197], we acknowledge the possibility for heterogeneous treatment effects when investigating responses of male and female characters revealing sexual orientation minority status. Third, we explore preference for a fictitious character rather than an actual human being. Factors influencing preferences for characters (e.g., prestige, popularity) might therefore play an important role. Our setting could be most analogous to the choice of supporting a prominent soccer player.
@@
We have dismissed the possibility that any actual or presumed change in the character’s performance is driving our results. This strongly suggests that the estimated cost of coming out is unlikely to be driven by factors other than stigmatization. This insight holds significant implications for policymakers aiming to develop interventions that effectively tackle stigmatization and improve the overall well-being of LGB individuals. Policies should be formulated to discourage discriminatory behavior, either by increasing its costs or by creating inclusive social environments that promote the acceptance of sexual minority individuals and reduce the stigma. Raising awareness about the reaction to sexual minority disclosure could be an important step to develop such a society. At the same time, policymakers may also consider providing resources and support to individuals who have recently come out, such as access to counseling and mental health services. By doing so, they can mitigate some of the negative outcomes that may arise from coming out.
[…]
[sorting = nty]
Appendix A Additional Results
@@
Figure A.1: Identities and contributions of characters in the donor pool for the Graves’ synthetic control displayed in Figure 4.1.
@@
Figure A.2: Identities and contributions of characters in the donor pool for the Graves’ synthetic controls displayed in Figure 4.2.
@@
Figure A.3: Shares of matches played in each role before and after the coming-out event. Players are divided into two groups based on their preferences for Graves before his disclosure.
@@
Figure A.4: Point estimates and simultaneous $95\%$ confidence bands allowing for clustering at the player level for the placebo $ATT\left(t\right)$ estimated under an unconditional parallel trend (left panels) and a conditional parallel trend assumption (right panels). Each row corresponds to a different version of the treatment.
Any Reduction Substantial Reduction Complete Abandonment (1) (2) (3) (4) (5) (6) $\overline{ATT\left(t\right)}$ 0.373 2.603 2.095 3.064 1.526 1.234 [-6.602, 7.347] [-6.855, 12.061] [-4.268, 8.457] [-4.037, 10.164] [-4.272, 7.324] [-4.951, 7.420] Conditional PT Players 1,038 1,038 1,038 1,038 1,038 1,038 Treated 857 857 738 738 586 586 Observations 89,650 89,650 89,650 89,650 89,650 89,650
Table A.1: Point estimates and $95\%$ confidence intervals for the average time ATT ($\overline{ATT\left(t\right)}$). Standard errors are clustered at the player level and computed using the multiplier bootstrap. Columns marked with checkmarks under “Conditional PT” display the results obtained with the doubly-robust approach. The remaining columns display the results obtained with the unconditional approach.
@@ Appendix B Robustness Checks
Figure B.1: Ratio of post- to pre-treatment RMSE for Graves and control characters. Only control characters with a pre-treatment RMSE of at least 1 are considered to exclude overfit specifications.
@@
Figure B.2: Robustness checks results. The upper panel shifts the coming-out event ten days earlier, with the new treatment date denoted by the vertical gray dashed line. The lower panel reports leave-one-out estimates of the synthetic control series, obtained by excluding one of the characters of Figure A.1 at a time from the donor pool.
@@ Appendix C Twisted Fate
In the League of Legends universe, Graves, described as a “gruff-looking, broad-shouldered, middle-aged man,” forms a criminal partnership with Twisted Fate, who is characterized as “a tall, handsome male with tanned skin, trimmed beard, and long dark hair.” Together, they engage in various illicit activities until, in the midst of a heist, Graves finds himself captured. This event leads Graves to perceive a betrayal by Twisted Fate, prompting a pursuit of revenge upon his escape. However, the duo ultimately decides to reconcile their differences and resumes their collaboration.343434 The complete background of Graves and Twisted Fate is available at https://leagueoflegends.fandom.com/wiki/Graves and https://leagueoflegends.fandom.com/wiki/Twisted_Fate.
@@
We investigate whether this implied revelation has captured the players’ attention in Figure C.1, illustrating the Google search interest for the queries “Twisted Fate gay” and “LoL Twisted Fate.” Throughout $2022$, we observe approximately no interest in the former query, with a small spike occurring during the week of the coming-out event, amounting to less than half of the spike associated with Graves. In contrast, the search interest for the latter query remains relatively steady over the year and is always lower than that for Graves, suggesting the greater popularity of Graves among players.353535 This is also suggested by Graves’s pick rates (see Section 3.1) being approximately 3 to 4 times higher than those of Twisted Fate, as displayed in Figure C.2. These results underscore the relatively low attention directed towards Twisted Fate from players, who were primarily focused on Graves and the explicit establishment of his sexual orientation. As a result, we concentrate our analysis on Graves and his disclosure for a more credible identification of the effects of coming out.
Figure C.1: Google search interest over time for the queries “Graves gay” and “Twisted Fate gay” (top panel) and the queries “LoL Graves” and “LoL Twisted Fate” (bottom panel). The dashed vertical line denotes the week of disclosure, and the shaded area highlights the League of Legends World Championship.
Figure C.2: Graves and Twisted Fate’s daily pick rates. The series are smoothed by a Nadaraya-Watson regression. The rainbow area highlights LGBT Pride Month.
@@ D.1 Simultaneous Treatments
As described in Section 2.2, the disclosure of Graves’ sexual orientation coincided with the start of LGBT Pride Month. This means that the coming-out event encompasses two treatments occurring at the same time, namely the announcement of Graves’ homosexuality and the introduction of visual and expressive elements in League of Legends that support the LGBT community.363636 See, e.g., [204] for a discussion on “simultaneous treatments” and methodologies for disentangling their effects under a Difference-in-Differences identification strategy.
@@
To explicitly account for the influence of the two treatments $C_{i}$ and $L_{i}$, we define the potential pick rates as $Y_{i,t}^{c,l}$. Then, for each period $t>T^{pre}$, the effect of the coming-out event on players’ preferences for Graves in (3.1) corresponds to:
$\tau_{t}=Y_{1,t}^{1,1}-Y_{1,t}^{0,0}$ (D.1)
Equation (D.1) shows why we need to be cautious in interpreting the estimated effects of Section 4.1 as solely stemming from the disclosure of Graves’ sexual orientation. Under an extended version of the SUTVA assumption (see Section D.2), we observe $Y_{1,t}=Y_{1,t}^{1,1}$ for all $t>T^{pre}$, and the estimator in (3.3) effectively targets the counterfactual series $Y_{1,t}^{0,0}$. Consequently, the estimated effects presented in Section 4.1 encompass the combined impacts of both disclosing Graves’ sexual orientation and his affiliation with the LGB community during LGBT Pride Month. This can be formalized as follows:
$\begin{split}\tau_{t}&=Y_{1,t}^{1,1}-Y_{1,t}^{0,0}\\ &=\underbrace{\left[Y_{1,t}^{1,1}-Y_{1,t}^{0,1}\right]}_{:=\tau_{t}^{C}}+\underbrace{\left[Y_{1,t}^{0,1}-Y_{1,t}^{0,0}\right]}_{:=\tau_{t}^{L}}\\ \end{split}$ (D.2)

### v3

@@ 5.5 Coming Out versus LGBT Pride Month
As described in Section 2.2, the disclosure of Graves’ sexual orientation coincided with the start of LGBT Pride Month. This means that the coming-out event encompasses two “simultaneous treatments” [[, see, e.g.,]]roller2023differences, namely the announcement of Graves’ homosexuality and the introduction of visual and expressive elements in League of Legends that support the LGBT community. It is therefore plausible that the findings presented in Section 4 may, to some extent, be influenced by the presence of LGBT Pride Month, which might elicit negative reactions from certain players, leading them to shift their preferences away from LGB characters. While this alternative perspective does not undermine the validity of our identification strategy, it does raise questions about our interpretation of the estimated effects as solely stemming from Graves’ disclosure.
@@
Figure 9: Composite LGB unit’s daily pick rates and synthetic control estimation results. The actual series is smoothed by a Nadaraya-Watson regression before applying the synthetic control estimator. The dashed vertical line denotes the day of disclosure, and the rainbow area highlights LGBT Pride Month.
We create a composite LGB unit by averaging the pick rates of Diana, Leona, Nami, and Neeko and employ the synthetic control estimator described in Section 3.2 to estimate the effect of LGBT Pride Month on players’ preferences for LGB characters. Then, under the assumption that the influence of LGBT Pride Month is uniform across all LGB characters, we can compare the results with those obtained for Graves to separate the impacts of coming out and LGBT Pride Month on players’ preferences for Graves. Intuitively, if the estimated impact of LGBT Pride Month on players’ preferences for LGB characters is small relative to the estimated impact of the coming-out event on players’ preferences for Graves, this suggests that the findings of Section 4 must be primarily attributed to Graves’ disclosure.
Figure 9 displays the actual and the synthetic pick rate series for the composite LGB unit. Overall, our analysis suggests that LGBT Pride Month had no impact on players’ preferences for LGB characters. Before the treatment, the synthetic control estimator closely aligns with the actual series, providing support for the estimator’s ability to predict the counterfactual series. After the treatment date, the synthetic control estimator continues to align with the actual series, confirming that the players’ preferences for LGB characters were unaffected by LGBT Pride Month. The average effect is estimated to be 0.215-0.215 percentage points (standard error: $3.052$), and the conventional $95\%$ confidence interval encompasses zero, indicating a failure to reject the null hypothesis of no effect. Under the homogeneity assumption discussed above, these findings support the interpretation that the estimated effects presented in Section 4 are primarily driven by Graves’ disclosure rather than being influenced by the broader context of LGBT Pride Month.
@@ 6 Conclusion
To bolster the credibility of stigma as the primary explanation for the estimated effects, we address and eliminate several alternative channels. First, we rule out the possibility that shifts in characters’ relative strengths could explain our estimated effect. Second, we show that players’ skills have no correlation with the choice to drop the character. Third, we provide evidence that switching to other characters does not affect the performance of the players involved and that the release of new characters in the post-treatment period is unlikely to serve as primary explanation behind the results. Fourth, we introduce a theoretical framework that formalizes the existence of two ”simultaneous treatments” and use information on other playable characters that belong to the LGB group to rule out that LGBT Pride Month serves as the main explanation for the observed result.
@@
We have dismissed the possibility that any actual or presumed change in the character’s strength is driving our results. This strongly suggests that the estimated cost of coming out is unlikely to be driven by factors other than stigmatization. This insight holds significant implications for policymakers aiming to develop interventions that effectively tackle stigmatization and improve the overall well-being of LGB individuals. Policies should be formulated to discourage discriminatory behavior, either by increasing its costs or by creating inclusive social environments that promote the acceptance of sexual minority individuals and reduce the stigma. Raising awareness about the reaction to sexual minority disclosure could be an important step to develop such a society, as previous research has demonstrated the value of discrimination awareness for minority groups outcomes [170]. At the same time, policymakers may also consider providing resources and support to individuals who have recently come out, such as access to counseling and mental health services. By doing so, they can mitigate some of the negative outcomes that may arise from coming out.
Appendix A Additional Figures and Tables
Figure A.I: Histograms of players’ total and average daily activity, measured in both number of matches played and hours spent in-game.
@@ Appendix A Additional Figures and Tables
Figure A.II: Identities and contributions of characters in the donor pool for the Graves’ synthetic control displayed in Figure 2.
Figure A.III: Ratio of post- to pre-treatment RMSE for Graves and control characters. Only control characters with a pre-treatment RMSE of at least 1 are considered to exclude overfit specifications.
Figure A.IV: Distribution of overall win rates and total matches played for players classified as above or below the median pre-treatment win rate.
@@
Figure A.V: Identities and contributions of characters in the donor pool for the Graves’ synthetic controls displayed in Figure 4.
@@
Figure A.VI: Shares of matches played in each role before and after the coming-out event. Players are divided into two groups based on their preferences for Graves before his disclosure.
Figure A.VII: Distribution of percentage reduction in average pick rates for Graves from pre-treatment to post-treatment among prior users.
@@
Figure A.VIII: Point estimates and simultaneous $95\%$ confidence bands allowing for clustering at the player level for the placebo $ATT(t)$. Each row corresponds to a different version of the treatment.
Figure A.IX: Bel’Veth’s daily pick rates. The series is smoothed by a Nadaraya-Watson regression. The rainbow area highlights LGBT Pride Month. The non-zero pick rate displayed in early January is a result of the smoothing.
Table A.I: Most popular characters.
Pre-treatment Post-treatment Top Jungle Mid Bottom Support Top Jungle Mid Bottom Support 1 Irelia LeeSin Ahri Jinx Karma Gangplank Viego Yone Ezreal Renata (12.007) (19.132) (13.62) (24.5) (19.522) (12.455) (19.419) (17.587) (29.02) (16.43) 2 Camille Graves Akali Ezreal Nautilus Fiora MonkeyKing Sylas Zeri Karma (10.4) (18.557) (12.816) (21.155) (15.302) (10.569) (15.064) (13.177) (19.716) (16.239) 3 Jayce Viego Yone Jhin Lulu Aatrox LeeSin Ahri Twitch Yuumi (9.756) (17.982) (12.197) (20.746) (14.322) (9.451) (12.78) (12.83) (14.751) (14.215) 4 Fiora Diana Yasuo Kaisa Nami Irelia Taliyah Yasuo Jhin Senna (8.777) (13.771) (10.611) (17.174) (11.781) (9.407) (12.322) (10.855) (14.531) (13.715) 5 Aatrox Khazix Viktor Lucian Pyke Kayle Graves Viktor Kaisa Lulu (8.616) (9.448) (10.525) (12.505) (11.401) (8.51) (12.319) (10.347) (13.317) (13.274)
Notes. Most popular characters by role based on average pick rates (displayed in parenthesis)
Table A.II: Regional results.
Synthetic Controls Regularized Synthetic Controls (1) (2) (3) (4) All characters Only non-substitutes All characters Only non-substitutes Panel 1: Europe $\hat{\tau} -$9.062 -11.719 -8.696 -10.670 $95\%$ CI [-15.832, -2.292] [-19.479, -3.958] [-14.210, -3.183] [-19.594, -1.758] N. Donors 5 5 6 8 RMSE 2.206 2.639 2.155 2.583 Pre-treatment average 14.618 14.618 14.618 14.618 Panel 2: Korea $\hat{\tau} -$10.015 -8.955 -10.299 -8.990 $95\%$ CI [-18.262, -1.768] [-20.502, 2.593] [-17.720, -2.878] [-18.368, 0.376] N. Donors 3 3 3 3 RMSE 6.211 6.550 6.249 6.594 Pre-treatment average 31.441 31.441 31.441 31.441 Panel 3: Latin America $\hat{\tau} -$6.212 -5.899 -5.206 -4.960 $95\%$ CI [-11.850, -0.574] [-12.449, 0.651] [-10.645, 0.233] [-12.531, 2.605] N. Donors 4 4 7 6 RMSE 2.575 2.552 2.514 2.522 Pre-treatment average 16.217 16.217 16.217 16.217 Panel 4: North America $\hat{\tau}$ 0.939 0.940 1.110 1.130 $95\%$ CI [ -4.583, 6.462] [ -6.259, 8.138] [ -3.412, 5.632] [ -5.807, 8.079] N. Donors 3 3 4 4 RMSE 4.669 4.669 4.681 4.685 Pre-treatment average 22.282 22.282 22.282 22.282
Notes. Point estimates and $95\%$ confidence intervals for $\hat{\tau}$. Additionally, the number of donors receiving a non-zero weight and the pre-treatment root mean squared error are displayed. Each panel reports the results obtained using only matches from a particular macro-region. Each column corresponds to a different specification, with the specifications differing solely in the employed estimator and donor pool composition.
@@ Appendix B Robustness Checks
This section presents a series of robustness checks that support the credibility of our findings. We begin with checks related to our main results, followed by additional analyses that validate the player heterogeneity findings.
B.1 Main Results
We examine the robustness of our main results from Section 4.1 to alternative estimation strategies and variations in the composition of the donor pool. In particular, we repeat our analysis employing a regularized synthetic control estimator (see Section 3.2) and explore different donor pool configurations focusing on characters from distinct roles. Notably, Graves is predominantly designed for and played in three of the possible roles within a team. Consequently, there is a possibility of spillover effects on other characters mainly played in these positions, as players transitioning away from Graves are likely to switch to these alternatives.404040 Table 1 and the findings of Section 5.2 support this intuition. To mitigate this potential for spillover effects, we restrict our donor pool to characters that are “non-substitutes” of Graves, that is, those primarily designed for the remaining two roles.414141 Graves is predominantly designed for and played in the top, jungle, and mid positions. Therefore, we consider characters designed for and played in the bottom and support positions as “non-substitutes.”
Table B.I: Main results.
Synthetic Controls Regularized Synthetic Controls (1) (2) (3) (4) All characters Only non-substitutes All characters Only non-substitutes $\hat{\tau} -$7.606 -7.161 -7.323 -7.080 $95\%$ CI [-13.542, -1.670] [-15.383, 1.062] [-11.471, -3.174] [-14.831, 0.665] N. Donors 4 4 4 5 RMSE 2.365 2.525 2.330 2.388 Pre-treatment average 18.558 18.558 18.558 18.558
Notes. Point estimates and $95\%$ confidence intervals for $\hat{\tau}$. Additionally, the number of donors receiving a non-zero weight and the pre-treatment root mean squared error are displayed. Each column corresponds to a different specification, with the specifications differing solely in the employed estimator and donor pool composition.
Table B.I displays the results. For any donor pool composition, the results are not sensitive to the choice of the regularization parameter, and point estimates are consistently negative. The results are consistent also quantitatively across all specifications, with a decline ranging between $40.98\%$ and $38.17\%$ of the pre-treatment average preferences for Graves. Overall, these results support our main finding of a substantial negative impact of the coming-out event on players’ preferences for Graves.
@@ B.1 Main Results
Figure B.I: Robustness checks for main results. The upper panel shifts the coming-out event ten days earlier, with the new treatment date denoted by the vertical gray dashed line. The lower panel reports leave-one-out estimates of the synthetic control series, obtained by excluding one of the characters of Figure A.II at a time from the donor pool.
We also assess the credibility of the synthetic control estimator by conducting a robustness check that artificially shifts the coming-out event ten days earlier. This backdating exercise allows us to evaluate the estimator’s predictive accuracy during a ten-day hold-out period [[, see e.g.,]]abadie2022synthetic. The upper panel of Figure B.I presents the results of this analysis. We observe three key findings. First, the estimated effects remain qualitatively and quantitatively consistent, confirming a negative and persistent impact of the coming-out event on players’ revealed preferences for Graves. Second, the synthetic control estimator demonstrates a good fit during the hold-out period, indicating its ability to accurately capture Graves’ behavior prior to the disclosure. Third, the actual and the synthetic series begin to diverge on the true day of disclosure, even when the estimator has no knowledge of the actual disclosure date. The absence of estimated effects before the coming-out event also lends support to the plausibility of a no-anticipation assumption [[, see e.g.,]]abadie2021using.
Finally, we conduct an additional robustness test by performing a leave-one-out exercise, where we repeatedly estimate the synthetic control series by excluding one character with non-zero estimated weights at a time from the donor pool [[, see e.g.,]]abadie2021using. The lower panel of Figure B.I presents the results of this analysis. Overall, our finding of a negative and persistent impact of the coming-out event on players’ preferences for Graves is robust to the exclusion of any particular character. Most of the leave-one-out synthetic series closely align with the main estimate, thus reinforcing the robustness of the main conclusion of our study. One leave-one-out series falls beneath the other synthetic series, suggesting a somewhat reduced, although still negative, impact. However, this series diverges from the actual series in the weeks prior to the treatment, which undermines the reliability of its results.424242 This series is obtained by excluding the character Ezreal from the donor pool. Figure A.II in Appendix A shows that this character receives the largest weight in our main specification. Therefore, the divergence of this leave-one-out series from the actual series is unsurprising.
B.2 Player Heterogeneity
We assess the robustness of our player heterogeneity results from Section 4.2 using the same backdating and leave-one-out exercises described in the previous subsection. Figure B.II displays the results. For both skill groups, we observe that even when the estimator has no knowledge of the actual disclosure date, the actual and synthetic series begin to diverge precisely on the true day of disclosure. The fit during the hold-out period is also strong, indicating good predictive performance. Moreover, most of the leave-one-out synthetic series closely track the main estimates, reinforcing the robustness of our findings.
(a) Players below median skill
(b) Players above median skill
Figure B.II: Robustness checks for player heterogeneity results. The upper panels shift the coming-out event ten days earlier, with the new treatment date denoted by the vertical gray dashed line. The lower panels report leave-one-out estimates of the synthetic control series, obtained by excluding one character with non-zero estimated weights at a time from the donor pool.
@@ Appendix C Twisted Fate
In the League of Legends universe, Graves, described as a “gruff-looking, broad-shouldered, middle-aged man,” forms a criminal partnership with Twisted Fate, who is characterized as “a tall, handsome male with tanned skin, trimmed beard, and long dark hair.” Together, they engage in various illicit activities until, in the midst of a heist, Graves finds himself captured. This event leads Graves to perceive a betrayal by Twisted Fate, prompting a pursuit of revenge upon his escape. However, the duo ultimately decides to reconcile their differences and resumes their collaboration.434343 The complete background of Graves and Twisted Fate is available at https://leagueoflegends.fandom.com/wiki/Graves and https://leagueoflegends.fandom.com/wiki/Twisted_Fate.
@@
We investigate whether this implied revelation has captured the players’ attention in Figure C.I, illustrating the Google search interest for the queries “Graves gay” and “Twisted Fate gay.” Throughout $2022$, we observe approximately no interest in the latter query, with a small spike occurring during the week of the coming-out event, amounting to less than half of the spike associated with Graves. Furthermore, the search interest for Twisted Fate is always lower than that for Graves, suggesting the greater popularity of Graves among players.444444 This is also suggested by Graves’s pick rates being approximately 3 to 4 times higher than those of Twisted Fate, as displayed in Figure C.II. These results underscore the relatively low attention directed towards Twisted Fate from players, who were primarily focused on Graves and the explicit establishment of his sexual orientation. As a result, we concentrate our analysis on Graves and his disclosure for a more credible identification of the effects of coming out.
Figure C.I: Google search interest over time for the queries “Graves gay” and “Twisted Fate gay.” The dashed vertical line denotes the week of disclosure, and the shaded area highlights the League of Legends World Championship.
Figure C.II: Graves and Twisted Fate’s daily pick rates. The series are smoothed by a Nadaraya-Watson regression. The rainbow area highlights LGBT Pride Month.
@@ D.1 Simultaneous Treatments
As described in Section 2.2, the disclosure of Graves’ sexual orientation coincided with the start of LGBT Pride Month. This means that the coming-out event encompasses two treatments occurring at the same time, namely the announcement of Graves’ homosexuality and the introduction of visual and expressive elements in League of Legends that support the LGBT community.454545 See, e.g., [197] for a discussion on “simultaneous treatments” and methodologies for disentangling their effects under a Difference-in-Differences identification strategy.
We recognize the potential influence of LGBT Pride Month on players’ preferences for characters by introducing the binary variable $L_{i}\in\left\{0,1\right\}$ to represent character $i$’s inclusion in the LGB community no later than $T^{pre}+1$. Consequently, we observe three distinct groups of units: the first group includes only Graves, with $C_{i}=L_{i}=1$; the second group includes only Diana, Leona, Nami, and Neeko, with $C_{i}=0$ and $L_{i}=1$; and the third group includes all other characters, with $C_{i}=L_{i}=0$.464646 Neglecting the presence of two simultaneous treatments and treating them as a single treatment does not invalidate the results of Section 4.1. It primarily affects their interpretation, which, without further investigation, could only be attributed to the combined effects of simultaneously receiving both treatments $C_{i}$ and $L_{i}$ - referred to as the coming-out event in the main body of the paper.
To explicitly account for the influence of the two treatments $C_{i}$ and $L_{i}$, we define the potential pick rates as $Y_{i,t}^{c,l}$. Then, for each period $t>T^{pre}$, the effect of the coming-out event on players’ preferences for Graves in (1) corresponds to:
$\tau_{t}=Y_{1,t}^{1,1}-Y_{1,t}^{0,0}$ (D.1)
Equation (D.1) shows why we need to be cautious in interpreting the estimated effects of Section 4.1 as solely stemming from the disclosure of Graves’ sexual orientation. Under an extended version of the SUTVA assumption (see Section D.2), we observe $Y_{1,t}=Y_{1,t}^{1,1}$ for all $t>T^{pre}$, and the estimator in (3) effectively targets the counterfactual series $Y_{1,t}^{0,0}$. Consequently, the estimated effects presented in Section 4.1 encompass the combined impacts of both disclosing Graves’ sexual orientation and his affiliation with the LGB community during LGBT Pride Month. This can be formalized as follows: